# Accelerating Large Language Model Pretraining via LFR Pedagogy: Learn, Focus, and Review

### v1

@@ A.2 Extended Analysis on Dropped and Retained Data Blocks
Figure 6: Cosine similarity heatmaps for dropped and retained data blocks during Phase 2 of pretraining for the GPT-2 124M (right) and 345M (left) models.
@@
In this section, we expand on the ablation study in Section 5.8 in order to better characterize the data blocks considered easy / hard.
Tables 12 and 14 provides examples of text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models respectively. Similarly, Tables 13 and 15 provide examples of data blocks retained by the models in Phases 2 and 4. We printed out and went over all the text dropped and retained in both Phases, and notice that text considered easy in phase 2 was more conversational, and those considered easy in phase 4 were more factual. This might indicate that the model first learned conversations and personal anecdotes, before being able to retain factual information. These findings are further corroborated by the examples of data retained in both phases. We are working on further analysis across different model families and sizes to strengthen this understanding.
Next, we continue the analysis of the cosine similarity heatmaps evaluated across training time and model parameter scales presented in Section 5.8. Here, we answer the following questions:
@@
    Are the data blocks dropped similar to those retained at any given training phase? Note that Section 5.8 presented this analysis only for Phase 4 of the 124M and 345M parameter models in Figure 4.
Figure 5 depicts the cosine similarity heatmap for the data blocks dropped by the 124M parameter model in Phase 4 (X-axis) with those dropped by the 345M parameter model in Phase 2 (Y-axis). The mean, standard deviation, and variance of the similarity are 0.43, 0.18, and 0.03 respectively. In contrast, the mean cosine similarity of data blocks dropped in Phase 2 of pretraining of both the models was 0.38 (Section 5.8 and Figure 3). This indicates that the smaller model "catches up" with the knowledge accumulated by the larger model, and considers similar block easy in Phase 4 as those considered easy by the larger model in Phase 2.
Next, in Figure 6, we plot the cosine similarity heatmap for the dropped and retained data blocks in Phase 2 for the 124M (left) and 345M (right) parameter models. The mean, variance, and standard deviations of the similarity are shown in Table 11. Observing the mean similarity value and heatmap in Table 10 and Figure 4, we find that the cosine similarity for dropped and retained data blocks is higher in Phase 2 than Phase 4 in case of the 345M parameter model. In contrast, the value remains high in both Phases for the 124M parameter model. This finding indicates that both the smaller and larger model start the training by being confused about similar data blocks. However, the larger capacity of the 345M parameter model allows it to learn the dataset well in Phases 2 and 3, and move on to more complex data blocks in Phase 4 (thus reducing the mean similarity in Phase 4). The smaller model continues remaining unsure about similar data blocks. Since we observed that the smaller model "catches up" with the training of the larger model (in Figure 5), we hypothesize that the smaller model will eventually display similar behaviour as the larger model once trained for longer iterations.

### v2

@@ A.5 Extended Analysis on Dropped and Retained Data Blocks for GPT-2
In this section, we expand on the ablation study in Section A.4 in order to better characterize the data blocks considered easy / hard.
Tables 14 and 16 provides examples of text blocks dropped in Phases 2 and 4 by the 345M and 124M parameter models respectively. Similarly, Tables 15 and 17 provide examples of data blocks retained by the models in Phases 2 and 4. We printed out and went over all the text dropped and retained in both Phases, and notice that text considered easy in phase 2 was more conversational, and those considered easy in phase 4 were more factual. This might indicate that the model first learned conversations and personal anecdotes, before being able to retain factual information. These findings are further corroborated by the examples of data retained in both phases. We are working on further analysis across different model families and sizes to strengthen this understanding.
Next, we continue the analysis of the cosine similarity heatmaps evaluated across training time and model parameter scales presented in Section A.4. Here, we answer the following questions:
@@
    Are the data blocks dropped similar to those retained at any given training phase? Note that Section A.4 presented this analysis only for Phase 4 of the 124M and 345M parameter models in Figure 9.
Figure 10 depicts the cosine similarity heatmap for the data blocks dropped by the 124M parameter model in Phase 4 (X-axis) with those dropped by the 345M parameter model in Phase 2 (Y-axis). The mean, standard deviation, and variance of the similarity are 0.43, 0.18, and 0.03 respectively. In contrast, the mean cosine similarity of data blocks dropped in Phase 2 of pretraining of both the models was 0.38 (Section A.4 and Figure 8). This indicates that the smaller model "catches up" with the knowledge accumulated by the larger model, and considers similar block easy in Phase 4 as those considered easy by the larger model in Phase 2.
Next, we plot the cosine similarity heatmap for the dropped and retained data blocks in Phase 2 for the 124M (left) and 345M (right) parameter models. The mean, variance, and standard deviations of the similarity are shown in Table 13. Observing the mean similarity value and heatmap in Table 12 and Figure 9, we find that the cosine similarity for dropped and retained data blocks is higher in Phase 2 than Phase 4 in case of the 345M parameter model. In contrast, the value remains high in both Phases for the 124M parameter model. This finding indicates that both the smaller and larger model start the training by being confused about similar data blocks. However, the larger capacity of the 345M parameter model allows it to learn the dataset well in Phases 2 and 3, and move on to more complex data blocks in Phase 4 (thus reducing the mean similarity in Phase 4). The smaller model continues remaining unsure about similar data blocks. Since we observed that the smaller model "catches up" with the training of the larger model (in Figure 10), we hypothesize that the smaller model will eventually display similar behaviour as the larger model once trained for longer iterations.